How many grams of NaN3 are required to produce 12.0 ft3 of N2 if the gas has a

          density of 1.25g/L?


1
Expert's answer
2021-03-22T06:34:58-0400

2NaN3 = 2Na + 3N2

12.0 ft3 = 340 l

n (N2) = V/22.4 = 340 / 22.4 = 15.2 mol

According to the equation, n (NaN3) = 2/3 x n (NaN2) = 2/3 x 15.2 = 10.1 mol

n = m/M

m = n x M

M (NaN3) = 65 g/mol

m (NaN3) = 10.1 x 65 = 657.7 g